ADM2341 Managerial Accounting Fall 2019

Solution: Chapter 6 Capstone Problems

Q1) The books of Lionel Company, wholesalers of hand held calculators, reflected the following revenues and expenses for various months during the year ended December 31, 2008.

August December May
Sales in units 24,000 28,000 22,000
Sales revenues $336,000 $392,000 $308,000
Cost of Goods Sold ($60,000) ($70,000) (55,000)
Gross Margin $276,000 $322,000 $253,000
Operating Expenses
Advertising expense (35,000) (35,000) (35,000)
Commissions (72,000) (84,000) (66,000)
Selling expense (16,600) (18,600) (15,600)
Operating income $152,400 $184,000 $136,400

1. Identify the above expenses as either variable, fixed, and mixed.

2. Separate each mixed expense into fixed and variable expense by using the high low method. State the cost formula for each expense.

3. Prepare a contribution income statement for the month of December.

Q2) Print Inc., of Ottawa, producers of finance and accounting books, reflected the following information at December 31, 2007, and 2008.

2007-12-31 2008-12-31
Production 100,000 units 200,000 units
Cost of Goods Manufactured $1,400,000 $2,600,000
Work in progress inventory, beginning $40,000 $20,000
Work in progress inventory, ending $20,000 $60,000
Direct materials inventory, per unit $4 $4
Direct labor, per unit $7.5 $7.5
Manufacturing overhead ? ?

Manufacturing overhead consists of both variable and fixed cost elements. Management wants to determine the overhead breakdown between variable and fixed cost per year.

Required:

1. For both years, determine total manufacturing costs. (Hint: Use a Cost of Goods Sold Manufacturing Schedule).

2. Determine the cost formula for manufacturing overhead by means of the high-low method of cost analysis. Express the variable portion in terms of variable cost per unit of production.

3. If 160,000 units are produced during a period what would be the total manufacturing costs?
Answer #1

1.if cost is fixed per unit it is a variable cost. if total cost is fixed in whole the cost is fixed cost. if cost is niether fixed per unit nor in total it is mixed cost.

august december may classification
cost of goods sold 2.5$[60000/24000] 2.5$[70000/28000] 2.5$[55000/22000] variable cost.
advertising 35000 35000 35000 fixed cost
commission 3[72000/24000 3[84000/28000 3[66000/22000 variable cost
selling expense 0.69[16600/24000 0.66[18600/28000 0.71[15600/22000 mixed cost

2.high low method

high low change
change in level 28000 22000 6000
change in cost 18600 15600 3000
variable cost = change in cost / change in level 0.5$ per unit

cost formula Y = a+bx

y total cost

a= fixed cost

x number of units

b = variable cost per unit

18600 = a + 28000*0.5

a=18600-14000

=4600$ is fixed cost

variable cost for 24000units = 24000*0.5=12000$

variable cost for 22000 units =0.5*22000=11000$

3.contribution margin statement for December

December
sales 392,000
less variable expense
cost of goods sold (70000)
commission expense (84000)
selling expense (14000)
Total variable expense (168000)
contribution margin 224,000 [392000-70000-84000-14000]
less fixed cost
selling cost (4600)
advertising expense (35000)
Total fixed cost (39600)
Net income(loss) $184,400 [224000-4600-35000]